[Hallb] [Clas_members] CLAS shifts this evening

orielly at jlab.org orielly at jlab.org
Tue Jul 27 08:48:50 EDT 2010

Hello Collaborators,

the target group got FROST down to 65 mK yesterday, below the goal of 70
mK. The first attempt to polarize last night unfortunately failed, with a
magnet quench, but the group is recooling FROST and is confident that they
will succeed in getting a polarized target today.

Shifts will restart TODAY at 16:00 (Tuesday, July 27)



Clas_members mailing list
Clas_members at jlab.org

More information about the Hallb mailing list